How much mass is converted ro energy per day in a nuclear power plant operated at a level of 100 MW?
1
Expert's answer
2021-07-02T13:44:03-0400

10 MW "= 10^{7} \frac {J}{sec}"

Energy produce per day "= 10^7\times 24 \times 60\times 60 =864 \times 10^{9}"

"E = \triangle m c^2"

"\triangle m = \frac {c^2}{E} = \frac{(3\times10^8)^2 }{864 \times 10^{9} } = 104.166\times 10^3 kg"